### v2

# Valentina Bais and Rafael Torres Scuola Internazionale Superiori di Studi Avanzati (SISSA) Via Bonomea 265 34136

Abstract: We introduce a simple cut-and-paste mechanism to construct both orientable and nonorientable four-manifolds from a given initial one. This mechanism alters the fundamental group while preserving other essential topological invariants. It avoids codimension two cut-and-paste fundamental group computations and fast tracks the search for fixed-point free involutions. The mechanism proves useful to unveil novel exotic irreducible smooth structures on closed four-manifolds with finite cyclic fundamental group, which include $\mathbb{Q}$-homology real projective four-spaces.
1. Introduction.
The study of existence of nondiffeomorphic smooth structures is a central theme in four-manifold theory, where the development of novel production techniques that pair well with a given diffeomorphism invariant to manufacture new examples is of pre-eminent importance. The main contribution of this paper is to record a simple construction mechanism and to sample its efficiency by building new irreducible smooth structures on closed four-manifolds. The procedure consists of four steps.
  1. I
    . Start with a pair $(M_{1},M_{2})$ of smooth four-manifolds that are homeomorphic, but not diffeomorphic. Remove the tubular neighborhood $\nu(\alpha_{i})$ of a simple loop $\alpha_{i}\subset M_{i}$ to produce a compact four-manifold
    $M_{i}^{0}\mathrel{\mathop{:}}=M_{i}\setminus\nu(\alpha_{i})$
    for $i=1,2$.
  2. II
    . Build a compact oriented four-manifold $S_{p}$ that is a $\mathbb{Q}$-homology $S^{1}\times D^{3}$ with boundary $\partial S_{p}=S^{1}\times S^{2}$ and a compact nonorientable four-manifold $N_{p}$ that is a $\mathbb{Q}$-homology $D^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ (the nonorientable three-disk bundle over the circle) with boundary $\partial N_{p}=S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ (the nonorientable two-sphere bundle over the circle). A detailed description of the compact four-manifolds $S_{p}$ and $N_{p}$ is given in Section 2.1 and Section 6.1, respectively. Both of these building blocks have the property that any self-diffeomorphism of their boundary extends to a self-diffeomorphism of the whole four-manifold; see Theorem 4 and Theorem 25.
    Theorem A.
    Let $p\in\mathbb{Z}_{>0}$ be odd.
    $\bullet$ For any self-diffeomorphism $h:S^{1}\times S^{2}\rightarrow S^{1}\times S^{2}$, there is a diffeomorphism $H_{p}:S_{p}\rightarrow S_{p}$ such that $H_{p}|_{\partial}=h$.
    $\bullet$ For any self-diffeomorphism $f:S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}\rightarrow S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$, there is a diffeomorphism $F_{p}:N_{p}\rightarrow N_{p}$ such that $F_{p}|_{\partial}=f$.
    Theorem A is a generalization of well-known foundational results of Laudenbach-Poénaru [28], César de Sá [9] and Miller-Naylor [30], and it might be of independent interest.
  3. III
    . For $i=1,2$, assemble new four-manifolds
    $M_{i}(p)\mathrel{\mathop{:}}=M_{i}^{0}\cup S_{p}$
    if $\alpha_{i}\subset M_{i}$ is an orientation-preserving loop or
    $M_{i}(p)\mathrel{\mathop{:}}=M_{i}^{0}\cup N_{p}$
    if it is an orientation-reversing one. This cut-and-paste procedure introduces fundamental group, in the sense that
    $\pi_{1}(M_{i}(p))\neq\pi_{1}(M_{i}),$
    and its codimension three nature avoids cumbersome fundamental group computations that are unavoidable in the codimension two case: see [5]. The property of the building blocks $S_{p}$ and $N_{p}$ mentioned in Theorem A has the following useful consequence.
    Theorem B.
    Let $M_{1}$ and $M_{2}$ be topological four-manifolds such that there are locally flat simple loops
    $\{\alpha_{M_{i}}\subset M_{i}:i=1,2\}$
    for which there is a homeomorphism
    $M_{1}\setminus\nu(\alpha_{M_{1}})\rightarrow M_{2}\setminus\nu(\alpha_{M_{2}}).$
    If $\alpha_{M_{i}}\subset M_{i}$ is an orientation-preserving loop, set $C_{p}\mathrel{\mathop{:}}=S_{p}$. Otherwise, let $C_{p}\mathrel{\mathop{:}}=N_{p}$. For $i=1,2$, define
    $M_{i}(p)\mathrel{\mathop{:}}=(M_{i}\setminus\nu(\alpha_{M_{i}}))\cup C_{p}.$
    The four-manifolds $M_{1}(p)$ and $M_{2}(p)$ are homeomorphic.
  4. IV
    . Find an invariant that distinguishes the diffeomorphism types of the new four-manifolds $M_{1}(p)$ and $M_{2}(p)$.
The crux to the efficiency of the procedure is its fourth step. In principle, one would hope that the diffeomorphism invariant that discerns $M_{1}$ from $M_{2}$ also tells apart the new four-manifolds $M_{1}(p)$ and $M_{2}(p)$. This is indeed the case in all the examples under consideration in this paper, where we reckoned the Heegaard Floer homology based invariants used by Levine-Lidman-Piccirillo in [29], the Ozsváth-Szabó invariants [31], the Seiberg-Witten invariants [20], [34, Section 2.4] (as employed by Stipsicz-Szabó [33, 34] and Baykur-Stipsicz-Szabó [8]), and the $\eta$-invariant as studied by Stolz [35].
As a warm up result, we paint a picture of the fourth step and the utility of the mechanism through an extension of a recent result of Levine-Lidman-Piccirillo. In [29], the authors constructed the first example of an inequivalent smooth structure on a closed smooth oriented four-manifold with definite intersection form: they built a four-manifold $\mathcal{R}$ that is homeomorphic but not diffeomorphic to the connected sum $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$, where $\Sigma_{2}$ is a $\mathbb{Q}$-homology four-sphere with $\pi_{1}(\Sigma_{2})=\mathbb{Z}/2$. Using the pair $(\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}},\mathcal{R})$ in the first step of our mechanism, we obtain the following generalization of their result.
Theorem C.
For every odd $p\in\mathbb{Z}_{>0}$, there is a closed smooth oriented four-manifold $\mathcal{R}_{2p}$ that is homeomorphic but not diffeomorphic to the connected sum $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$, where $\Sigma_{2p}$ is a $\mathbb{Q}$-homology four-sphere with cyclic fundamental group of order $2p$. The diffeomorphism types are distinsguished by the Ozsváth-Szabó invariant.
To distinguish the smooth structures in [29, Theorem 1.2], Levine-Lidman-Piccirillo show that the universal covers ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{E}$ of the four-manifolds $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{R}$ are non-diffeomorphic by means of their Ozsváth-Szabó invariant [31]. For every odd $p\in\mathbb{Z}_{>0}$, we show that the degree-two coverings of our four-manifolds are obtained from ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{E}$, respectively, by taking a connected sum with a $\mathbb{Q}$-homology four-sphere. This observation, together with the properties of the Heegaard Floer smooth four-dimensional invariant, allows us to canonically adapt the proof of [29, Theorem 1.2] to distinguish such coverings and, hence, also the four-manifolds $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ and $\mathcal{R}_{2p}$.
Our next theorem was inspired by recent work of Stipsicz-Szabó [33, 34] and Baykur-Stipsicz-Szabó [8], where they built infinitely many nondiffeomorphic irreducible smooth structures on $\Sigma_{4m+2}\mathbin{\#}\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ for $\Sigma_{4m+2}$ a $\mathbb{Q}$-homology four-sphere with fundamental group $\mathbb{Z}/(4m+2)$. Recall that a smooth four-manifold $M$ is said to be irreducible if for every smooth connected sum decomposition $M=M_{1}\mathbin{\#}M_{2}$ either $M_{1}$ or $M_{2}$ is a homotopy four-sphere [20, Definition 10.1.17].
Theorem D.
Let $(n,p)\in\mathbb{Z}_{>0}\times\mathbb{Z}_{\geq 3}$ and assume $p\geq 3$ is odd. Let $\Sigma_{q,k}$ be a $\mathbb{Q}$-homology four-sphere with finite cyclic fundamental group of order $q$ and Kirby-Siebenmann invariant $\operatorname{KS}(\Sigma_{q,k})=k$. Let $\widetilde{X}_{n}$ be a closed smooth oriented simply connected four-manifold that satisfies the following properties.
  • There is an orientation-preserving smooth fixed-point-free involution
    (1.1) $\tau_{n}:\widetilde{X}_{n}\rightarrow\widetilde{X}_{n}$
    and the quotient four-manifold $\widetilde{X}_{n}/\tau_{n}$ is homeomorphic to a connected sum $Y\mathbin{\#}\Sigma_{2,k}$, where $Y$ is a closed simply connected topological four-manifold with Kirby-Siebenmann invariant given by $\operatorname{KS}(Y)=k=\operatorname{KS}(\Sigma_{2p,k})$.
  • The four-manifold $\widetilde{X}_{n}$ has exactly two Seiberg-Witten basic classes ±$\pm K_{n}\in\mathcal{B}_{\widetilde{X}_{n}}$ and $SW_{\widetilde{X}_{n}}(\pm K_{n})=\pm n^{2}$.
Then, there exists a closed smooth oriented irreducible four-manifold $X_{n}(2p)$ with finite fundamental group of order $2p$ that satisfies the following properties.
  • There is a homeomorphism
    (1.2) $X_{n}(2p)\rightarrow Y\mathbin{\#}\Sigma_{2p,k}.$
  • The double cover $\widehat{X}_{n}(p)\rightarrow X_{n}(2p)$ corresponding to the index-two subgroup $\mathbb{Z}/p\subset\pi_{1}(X_{n}(2p))=\mathbb{Z}/2p$ decomposes as
    (1.3) $\widehat{X}_{n}(p)=\widetilde{X}_{n}\mathbin{\#}\Sigma_{p}.$
  • There are infinitely many pairwise nondiffeomorphic irreducible four-manifolds in the homeomorphism class of $Y\mathbin{\#}\Sigma_{2p,k}$ and these diffeomorphisms types are distinguished by the Seiberg-Witten invariant of their double covers.
If the four-manifold $\widetilde{X}_{n}$ does admit an involution (1.1), but has more than two Seiberg-Witten basic classes, assume it is minimal. Then, there exists a closed smooth oriented irreducible four-manifold $X_{n}(2p)$ with finite fundamental group of order $2p$ for which there is a homeomorphism (1.2) whose double cover is (1.3).
In [33, 34, 8], the Seiberg-Witten invariants are used to distinguish the diffeomorphism types of their examples in the following way. Notice that a diffeomorphism $h:M_{1}\rightarrow M_{2}$ between two closed oriented four-manifolds lifts to a diffeomorphism of the degree-two coverings $\hat{h}:\widehat{M}_{1}\rightarrow\widehat{M}_{2}$, whenever the index-two subgroups of $\pi_{1}(M_{i})$ identifying $\widehat{M}_{i}\rightarrow M_{i}$ correspond under the map induced by $h$; see Lemma 9. Stipsicz-Szabó obstruct the existence of the diffeomorphism $h$ through the diffeomorphism invariance of the Seiberg-Witten function $\operatorname{SW}_{\widehat{M}_{i}}:\operatorname{Spin}^{\mathbb{C}}(\widehat{M}_{i})\rightarrow\mathbb{Z}$ for $i=1,2$. Moreover, they conclude on the irreducibility of $\widehat{M}_{i}$ and $M_{i}$ through their Seiberg-Witten basic classes. We follow the same approach in the proof of Theorem D. In our case, however, the degree-two coverings of the four-manifolds of Theorem D do decompose smoothly as a connected sum $M\mathbin{\#}\Sigma_{p}$. In the proof of Theorem D, we couple the computations of the Seiberg-Witten invariants of Stipsicz-Szabó with a result of Kotschick-Morgan-Taubes [27, Proposition 2] that states that there are $\text{Spin}^{\mathbb{C}}$-structures on $M$ and $M\mathbin{\#}\Sigma_{p}$ for which the values of their Seiberg-Witten invariants coincide; see Section 4.1.
As a sample of the efficiency of our simple procedure, we construct new examples of irreducible smooth structures in the following theorem.
Theorem E.
Let $(m,n,p)\in\mathbb{Z}_{>0}\times\mathbb{Z}_{>0}\times\mathbb{Z}_{>0}$ be positive integers that satisfy $4+5m\geq n$, $4+5n\geq m$, $m=0\mod 2$ and $p\neq 0\mod 2$. If $m=n$, suppose $n>7$. There is a closed smooth irreducible four-manifold $X_{m,n}$ with finite cyclic fundamental group of order $2p$ that is homeomorphic to the connected sum
$m{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\#n\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}\#\Sigma_{2p}.$
The integers in the statement of Theorem E decompose the second Betti number of $X_{m,n}$ into
$b_{2}^{+}(X_{m,n}(2p))=m$ and $b_{2}^{-}(X_{m,n}(2p))=n$.
The case of $m\neq 0\mod 2$ and $m\leq n$ of Theorem E is well-known to hold for every finite cyclic fundamental group and for infinitely many pairwise nondiffeomorphic irreducible smooth structures by modifying the constructions of closed simply connected irreducible four-manifolds of Akhmedov-Baldridge-Baykur-Kirk-Park [2], Akhmedov-Park [3], Baldridge-Kirk [5, 6], Baykur-Hamada [7], Fintushel-Stern [14] and Gompf [18] as already done in [3, 6, 8, 37, 38]. The case $m=0\mod 2$ and $p=1$ is due to Stipsicz-Szabó [33, 34], Baykur-Stipsicz-Szabó [8] and Arabadji-Morgan [4]. Our proof of Theorem E is based on coupling their results with Theorem D, and it is presented in Section 5.
Concerning the nonorientable realm, we construct new smooth structures on $\mathbb{Q}$-homology real projective four-spaces and non-smoothable copies of these four-manifolds through our mechanism, and extend results of Cappell-Shaneson [10], Fintushel-Stern [12] and Ruberman [32]. In order to build the novel four-manifolds, we start with Cappell-Shaneson’s exotic $\mathbb{R}P^{4}$ or the non-smoothable four-manifold $\ast\mathbb{R}P^{4}$ that is homotopy equivalent to the real projective four-space. We address the fourth step of our construction mechanism by computing the $\eta$-invariant [35] of the smooth four-manifolds involved in Section 6.4. Our main result in this case is the following theorem.
Theorem F.
For every odd $p\in\mathbb{Z}_{>0}$, there are closed smooth irreducible nonorientable four-manifolds $A(2p)$ and $B(2p)$ that satisfy the following properties.
  1. (1)
    The Euler characteristics of these four-manifolds are equal to
    $\chi(A(2p))=1=\chi(B(2p))$
    and their fundamental groups are isomorphic to
    $\pi_{1}(A(2p))=\mathbb{Z}/2p=\pi_{1}(B(2p)).$
  2. (2)
    There is a homeomorphism $A(2p)\rightarrow B(2p)$.
  3. (3)
    The orientation 2-cover $\widehat{A}(2p)$ is diffeomorphic to the orientation 2-cover $\widehat{B}(2p)$.
  4. (4)
    The universal covers $\widetilde{A}(2p)$ and $\widetilde{B}(2p)$ are diffeomorphic to the connected sum $(2p-1)(S^{2}\times S^{2})$.
  5. (5)
    There is no diffeomorphism
    $A(2p)\mathbin{\#}(k-1)(S^{2}\times S^{2})\rightarrow B(2p)\mathbin{\#}(k-1)(S^{2}\times S^{2})$
    for any $k\in\mathbb{Z}_{>0}$.
  6. (6)
    The connected sums $A(2p)\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}$ and $B(2p)\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}$ are diffeomorphic.
Moreover, there is a non-smoothable closed topological four-manifold $\ast A(2p)$ that is homotopy equivalent to $A(2p)$.
We have collected quintessential properties of the known inequivalent smooth structures on nonorientable four-manifolds in the statement of Theorem F for the convenience of the reader.
The paper is organized as follows. Section 2 contains a detailed description of the building block $S_{p}$ and the construction mechanism in the orientable case as well as their basic properties. This includes the proofs of Theorem A and Theorem B in the oriented case. A proof of Theorem C is found in Section 3, while Section 4 accommodates a proof of Theorem D. The proof of Theorem E is discussed in Section 5. Section 6 is devoted to the study of the construction mechanism for nonorientable four-manifolds. In Section 6.1, we give a detailed description of the building block $N_{p}$ and the construction mechanism in the nonorientable case along with their basic properties. This includes the proofs of Theorem A and Theorem B in the nonorientable case. Theorem F is proven in Section 6.5.
We follow the following conventions and notation in the sequel. All our maps and manifolds are smooth unless it is stated otherwise. A pair $(X,\mathfrak{t})$ given by a four-manifold $X$ with a fixed $\operatorname{Spin}^{\mathbb{C}}$-structure $\mathfrak{t}$ will be called a $\operatorname{Spin}^{\mathbb{C}}$-structure.
1.1. Acknowledgements
V. B. is a member of GNSAGA – Istituto Nazionale di Alta Matematica ‘Francesco Severi’, Italy. R. T. thanks Stefan Bauer, Bob Gompf, Paul Kirk, Tye Lidman, Porter Morgan, Tom Mrowka, Mark Powell, Cliff Taubes, Mike Usher and Zoltan Szabó for useful e-mail exchanges.
2. Constructions of oriented four-manifolds.
2.1. Cut-and-paste procedure: oriented setting
We construct the compact oriented four-manifold $S_{p}$ of the second step of our procedure as follows. Let $\alpha\subset S^{1}\times S^{3}$ be a simple loop whose homotopy class generates the infinite cyclic group $\pi_{1}(S^{1}\times S^{3})=\mathbb{Z}$. Denote by $\alpha_{p}\subset S^{1}\times S^{3}$ the simple loop that represents the homology class $[\alpha]^{p}\in H_{1}(S^{1}\times S^{3};\mathbb{Z})=\mathbb{Z}$. The tubular neighborhood of this codimension three submanifold is $\nu(\alpha_{p})=S^{1}\times D^{3}$. The main building block for our cut-and-paste construction in the orientable setting is the compact four-manifold
(2.1) $S_{p}=(S^{1}\times S^{3})\setminus\nu(\alpha_{p})$
with boundary $\partial S_{p}=S^{1}\times S^{2}$.
Example 1.
The closed smooth oriented four-manifold
(2.2) $\Sigma_{p}=(D^{2}\times S^{2})\cup S_{p}$
is a $\mathbb{Q}$-homology four-sphere whose fundamental group is finite cyclic of order $p\in\mathbb{Z}_{>0}$, i.e. $\pi_{1}(\Sigma_{p})=\mathbb{Z}/p$.
A description of the cyclic coverings of the building block $S_{p}$ is given in our next lemma, which is of great utility in the sequel.
Lemma 2.
Let $S_{p}$ be the four-manifold with boundary defined in (2.1).
$\bullet$ For every $p\in\mathbb{Z}_{>0}$, there is a double covering
$D_{p}\xrightarrow{2:1}S_{p}$
that corresponds to the index-two subgroup $2\mathbb{Z}\subset\mathbb{Z}$.
$\bullet$ If $p$ is odd, then there is a diffeomorphism $D_{p}\rightarrow S_{p}$.
$\bullet$ If $p=2q$, then there is a diffeomorphism
$D_{p}\rightarrow(S^{1}\times S^{3})\setminus\nu(\alpha_{q}\sqcup\alpha_{q}^{\prime})$
where $\alpha_{q}^{\prime}$ is a parallel copy of $\alpha_{q}$.
[Proof]For any $p\in\mathbb{Z}_{>0}$, the universal cover of $S_{p}$ is
$\widetilde{S}_{p}=\mathbb{R}\times(S^{3}\setminus\{B_{1},\dots,B_{p}\})$
where $\{B_{1},\dots,B_{p}\}\subset S^{3}$ are disjointly embedded smooth $3$-balls. The $\mathbb{Z}$-action on $\widetilde{S}_{p}$ is generated by the map
$\phi:\widetilde{S}_{p}\rightarrow\widetilde{S}_{p}$
$(t,x)\mapsto(t+1,f_{p}(x))$
where $f_{p}$ is the restriction of a smooth automorphism of $S^{3}$ which cyclically permutes the $3$-balls $B_{1},\dots,B_{p}$. The degree-two covering of $S_{p}$ associated to the subgroup $2\mathbb{Z}\subset\mathbb{Z}=\pi_{1}(S_{p})$ is given by the quotient of $\widetilde{S}_{p}$ by the $\mathbb{Z}$-action generated by the map $\phi^{2}$. The conclusion follows from the fact that the square of a cyclic permutation of odd length is still cyclic, while the square of a cyclic permutation of even length factors as a product of two cycles of half the original length.
Remark 3.
In the following, it will be useful to consider the genus one Heegaard splitting of the 3-sphere
(2.3) $S^{3}=(S^{1}\times D^{2})\cup(D^{2}\times S^{1}).$
We can assume that the map $f_{p}$ in the proof of Lemma 2 is a rotation of $S^{3}$ of angle $2\pi/p$, which restricts to a rotation along the $D^{2}$-factor on the solid torus on the left side of (2.3) and to a rotation along the $S^{1}$-factor on the solid torus on the right side. Moreover, we can choose the three-balls $B_{1},\dots,B_{p}$ to be symmetrically embedded in the solid torus on the right side of (2.3), where they are permuted by the action of $f_{p}$.
We now prove Theorem A in the orientable case. As it has been already mentioned, this theorem is an extension of a foundational result of Laudenbach-Poenarú [28], which states that any self-diffeomorphism of $S^{1}\times S^{2}$ extends to a self-diffeomorphism of $S^{1}\times D^{3}$.
Theorem 4.
(cf. Theorem A). For every odd integer $p\geq 1$ and for any self-diffeomorphism
$h:S^{1}\times S^{2}\rightarrow S^{1}\times S^{2},$
there is a diffeomorphism
$H_{p}:S_{p}\rightarrow S_{p}$
such that $H_{p}|_{\partial}=h$.
[Proof]A result of Gluck [17] states that the diffeotopy group of $S^{1}\times S^{2}$ is isomorphic to $\mathbb{Z}/2$ and it is generated by the self-diffeomorphism
(2.4) $\varphi:S^{1}\times S^{2}\rightarrow S^{1}\times S^{2}$
given by
$(\theta,x)\mapsto(\theta,r_{\theta}(x)),$
where $r_{\theta}$ is the rotation of $S^{2}$ of angle $\theta$ around a fixed axis [17], [20, p. 156]. Such map is known as a Gluck twist. In order to prove Theorem 4, it is enough to show that the diffeomorphism $\varphi$ admits an extension $\Phi_{p}:S_{p}\rightarrow S_{p}$ for any odd $p\in\mathbb{Z}_{>0}$. We explicitly build the diffeomorphism $\Phi_{p}$ by using the identification
$S_{p}\cong(\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i}))/\sim_{\phi}$
coming from Lemma 2 as follows. Without loss of generality, we work in the setting of Remark 3. In particular, the deck transformation $\phi$ and the collection of three-balls $B_{1},\dots,B_{p}\subset D^{2}\times S^{1}$ are defined as in the proof of Lemma 2.
Consider the map
$F:\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})\rightarrow S^{3}\setminus\sqcup_{i=1}^{p}B_{i}$
defined by
$(t,x)\mapsto r_{t}(x),$
where $r_{t}$ denotes the rotation of $S^{3}$ of an angle $2\pi t$ that restricts to a rotation along the $D^{2}$-factor on the solid torus $D^{2}\times S^{1}$ of the Heegaard decomposition of $S^{3}$ in Remark 3. Such solid torus contains the three-balls $B_{1},\dots,B_{p}$. Notice that for every $t\in[0,1]$, the map $F(t,\cdot)$ is an $f_{p}$-equivariant self-diffeomorphism of $S^{3}\setminus\sqcup_{i=1}^{p}B_{i}$, where $f_{p}$ can be chosen to act on the genus one Heegaard splitting of $S^{3}$ described in Remark 3. We can now define the automorphism
$\Phi:\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})\rightarrow\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})$
$(t,x)\mapsto(t,F(t,x)).$
It is immediate to check that $\Phi$ is $\phi$-equivariant and, thus, it induces a self-diffeomorphism $\Phi_{p}:S_{p}\rightarrow S_{p}$. In particular, the restriction of $\Phi_{p}$ to $\partial S_{p}=S^{1}\times S^{2}$ is a composition of $p$ Gluck twists, which is again a Gluck twist since $p$ is odd. This concludes the proof.
As mentioned in the first step of our introduction, remove the tubular neighborhood of a framed simple loop $\alpha_{M}\subset M$ from a closed oriented four-manifold $M$ to obtain a compact oriented four-manifold $M\setminus\nu(\alpha_{M})$ with boundary $\partial(M\setminus\nu(\alpha_{M}))=S^{1}\times S^{2}$. Cap off its boundary with the compact four-manifold defined in (2.1) to construct a closed oriented four-manifold
(2.5) $M(p)\mathrel{\mathop{:}}=(M\setminus\nu(\alpha_{M}))\cup S_{p}.$
Theorem 4 has the orientable case of Theorem B as a consequence, which guarantees that the third step described in our introduction goes through.
Corollary 5.
(cf. Theorem B). Let $M_{1}$ and $M_{2}$ be topological four-manifolds and let $\alpha_{M_{i}}\subset M_{i}$ be a locally flat orientation-preserving simple framed loop for $i=1,2$. Suppose there is a homeomorphism
$f:M_{1}\setminus\nu(\alpha_{M_{1}})\rightarrow M_{2}\setminus\nu(\alpha_{M_{2}}).$
The four-manifolds
$M_{1}(p):=(M_{1}\setminus\nu(\alpha_{M_{1}}))\cup S_{p}\quad\text{and}\quad M_{2}(p)\mathrel{\mathop{:}}=(M_{2}\setminus\nu(\alpha_{M_{2}}))\cup S_{p}$
are homeomorphic.
The reader is directed to [16, Chapter 5] regarding existence and uniqueness results of the tubular neighborhood of a locally flat simple loop in a topological four-manifold.
Independently of Corollary 5, the simplicity of the construction makes it straight-forward to compute several invariants of the new four-manifold (2.5) as we now sample.
Lemma 6.
The closed oriented four-manifold $M(p)$ defined in (2.5) has the following basic topological invariants for every $p\in\mathbb{Z}_{>0}$.
  1. (1)
    The Euler characteristic is $\chi(M(p))=\chi(M)$.
  2. (2)
    The signature is $\sigma(M(p))=\sigma(M)$.
  3. (3)
    The equivariant intersection forms $\widetilde{I}_{M}$ and $\widetilde{I}_{M(p)}$ are isomorphic.
  4. (4)
    The second Stiefel-Whitney class is $w_{2}(M(p))=w_{2}(M)$.
  5. (5)
    If the homotopy class of the simple loop $\alpha_{M}\subset M$ generates the cyclic fundamental group $\pi_{1}(M)=\mathbb{Z}/q$, then the fundamental group of $M(p)$ is $\pi_{1}(M(p))=\langle t:t^{q\cdot p}=1\rangle$.
[Proof]We provide a proof of the third item. Let $\alpha_{M}\subset M$ be a simple loop as in (2.5) and denote by $\widetilde{\alpha}_{M}\subset\widetilde{M}$ its preimage in the universal cover $\widetilde{M}$ of $M$. The universal cover $\widetilde{M}(p)$ of $M(p)$ is constructed by gluing together copies of $\widetilde{M}\setminus\nu(\widetilde{\alpha}_{M})$ and $\widetilde{S}_{p}/\sim_{\phi^{k}}$ by matching each boundary component of $\widetilde{M}\setminus\nu(\widetilde{\alpha}_{M})$ with a boundary component of $\widetilde{S}_{p}/\sim_{\phi^{k}}$, where $\widetilde{S}_{p}$ is the universal covering of $S_{p}$ (see Lemma 2) and $k$ is the order of $[\alpha_{M}]\in\pi_{1}(M)$. In particular, if $k=\infty$ we set $\widetilde{S}_{p}/\sim_{\phi^{\infty}}=\widetilde{S}_{p}$. It follows that the equivariant intersection form $\widetilde{I}_{M(p)}$ of $M(p)$ is
$\widetilde{I}_{M(p)}=\widetilde{I}_{M}\otimes_{\mathbb{Z}[\pi_{1}(M)]}\mathbb{Z}[\pi_{1}(M(p))].$
A proof of the other items of Lemma 6 is obtained from straight-forward homological computations, Novikov’s additivity and the Seifert-van Kampen theorem.
Lemma 7.
Let $M$ be a closed oriented four-manifold and let $\alpha_{M}\subset M$ be a simple framed loop whose homotopy class is $[\alpha_{M}]=0\in\pi_{1}(M)$. There is a diffeomorphism
(2.6) $M(p)=(M\setminus\nu(\alpha_{M}))\cup S_{p}\rightarrow M\mathbin{\#}\Sigma_{p},$
where $\Sigma_{p}$ is a $\mathbb{Q}$-homology four-sphere with finite cyclic fundamental group of order $p$.
[Proof]The existence of the diffeomorphism (2.6) can be seen from the decompositions (2.2) in Example 1 and
(2.7) $M=M\mathbin{\#}S^{4}=M\mathbin{\#}((D^{2}\times S^{2})\cup(S^{1}\times D^{3})).$
Lemma 2 and Lemma 7 have the following consequence.
Lemma 8.
Let $M(2)$ be a closed oriented four-manifold whose universal cover is $\widetilde{M}$ and let $\alpha\subset M(2)$ be a closed simple framed loop such that its homotopy class generates the group $\pi_{1}(M(2))=\mathbb{Z}/2$. If $p\in\mathbb{Z}_{>0}$ is odd, then the degree-two cover $\widehat{M}(p)$ of the closed oriented four-manifold
$M(2p)=(M(2)\setminus\nu(\alpha))\cup S_{p}$
is diffeomorphic to
$\widehat{M}(p)=\widetilde{M}\mathbin{\#}\Sigma_{p}.$
We finish this section with the following observation regarding lifts of diffeomorphisms to coverings, whose proof is left to the reader.
Lemma 9.
Let
(2.8) $f:M_{1}\rightarrow M_{2}$
be a diffeomorphism between two four-manifolds $M_{1}$ and $M_{2}$ with fundamental group $\pi_{1}(M_{i})=\mathbb{Z}/2p$ and let $\widetilde{M}_{i}$ be the degree-two covering of $M_{i}$ for $i=1,2$. The diffeomorphism (2.8) lifts to a diffeomorphism
$\tilde{f}:\widetilde{M}_{1}\rightarrow\widetilde{M}_{2}.$
Although Lemma 9 clearly holds under much less restrictive hypothesis, this phrasing suffices for the purposes in this manuscript.
3. Examples of Theorem C.
In order to construct the four-manifolds $\mathcal{R}_{2p}$ of Theorem C, we make heavy use of the tools developed by Levine-Lidman-Piccirillo in [29].
3.1. Raw materials for Theorem C
We will employ as a raw material the exotic four-manifolds $\mathcal{E}$ and $\mathcal{R}$ constructed in [29, Theorem 1.2].
Theorem 10 (Levine-Lidman-Piccirillo [29]).
  1. (1)
    There exists a smooth four-manifold $\mathcal{E}$, which is homeomorphic but not diffeomorphic to ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$.
  2. (2)
    There exists a smooth four-manifold $\mathcal{R}$, which is homeomorphic but not diffeomorphic to $\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$, where $\Sigma_{2}$ is a $\mathbb{Q}$-homology four-sphere with $\pi_{1}(\Sigma_{2})\cong\mathbb{Z}/2$.
We refer the reader to [29, Section 6.2] for the precise definition of the four-manifolds $\mathcal{E}$ and $\mathcal{R}$. The four-manifolds of Theorem C are defined by applying the construction described in Section 2.1 to the exotic four-manifold $\mathcal{R}$ in the second item of Theorem 10.
For any integer $p\geq 1$, we define the closed oriented smooth four-manifold
(3.1) $\mathcal{R}_{2p}=(\mathcal{R}\setminus\nu(\alpha_{\mathcal{R}}))\cup S_{p}$
where $\alpha_{\mathcal{R}}\subset\mathcal{R}$ is a simple closed curve whose homotopy class generates the group $\pi_{1}(\mathcal{R})=\mathbb{Z}/2$. The homeomorphism type of each four-manifold (3.1) is pinned down in the following lemma.
Lemma 11.
For every integer $p\geq 1$, there is a homeomorphism
(3.2) $\mathcal{R}_{2p}\rightarrow\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}},$
where $\Sigma_{2p}$ is a $\mathbb{Q}$-homology four-sphere with fundamental group $\mathbb{Z}/{2p}$.
[Proof]The $\mathbb{Q}$-homology four-sphere on the right side of (3.2) can be deconstructed as
(3.3) $\Sigma_{2p}=(\Sigma_{2}\setminus\nu(\alpha_{\Sigma_{2}}))\cup S_{p},$
where $\alpha_{\Sigma_{2}}\subset\Sigma_{2}$ is a simple loop whose homotopy class generates the fundamental group $\pi_{1}(\Sigma_{2})\cong\mathbb{Z}/2\cong\pi_{1}(\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})$. The second clause of Theorem 10 implies that there is a homeomorphism
$(\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})\setminus\nu(\alpha_{\Sigma_{2}})\rightarrow\mathcal{R}\setminus\nu(\alpha_{\mathcal{R}}).$
Such homeomorphism can be extended to a homeomorphism as in (3.2) by the decomposition of $\Sigma_{2p}$ in (3.3) and Corollary 5.
Remark 12.
Alternatively, the existence of a homeomorphism as in Lemma 11 follows from work of Hambleton-Kreck [21, Theorem C]; cf. [34, Theorem 2.1].
To distinguish the diffeomorphism class of $\mathcal{R}_{2p}$ from the one of the connected sum $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$, we look at the double covering
$\mathcal{E}_{p}\xrightarrow{2:1}\mathcal{R}_{2p}$
corresponding to the index-two subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p=\pi_{1}(\mathcal{R}_{2p})$.
Proposition 13.
If $p\in\mathbb{Z}_{>0}$ is odd, the degree-two covering of $\mathcal{R}_{2p}$ corresponding to the index-two subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p$ is diffeomorphic to the four-manifold
$\mathcal{E}_{p}=(\mathcal{E}\setminus\nu(\gamma))\cup S_{p}=\mathcal{E}\mathbin{\#}\Sigma_{p},$
where $\mathcal{E}$ is the exotic four-manifold of the first clause of Theorem 10, $\gamma\subset\mathcal{E}$ is a simple loop and $S_{p}$ is the four-manifold defined in (2.1).
[Proof]This follows from the definition of $\mathcal{R}$ in [29] as the quotient of $\mathcal{E}$ by a fixed-point-free orientation-reversing involution, together with Lemma 8.
3.2. Diffeomorphism obstruction
In light of the definition of $\mathcal{E}$ in [29, Section 6.2], Proposition 13 implies the existence of a decomposition
(3.4) $\mathcal{E}_{p}=V\cup_{\zeta}V_{p},$
where $V=C_{0}\cup_{\phi_{0}\circ\sigma}Z$ is the building block defined in [29, Construction 6.4]. In particular, $C_{0}$ is a contractible four-manifold, $Z$ consists of five $-1$-framed $2$-handles attached to $\partial C_{0}$, $\partial V$ is the result of 00-surgery on $S^{3}$ along the Stevedore knot and $\zeta$ is any of its orientation-reversing fixed-point-free involutions. On the other hand, for every $p\geq 1$, $V_{p}$ is defined to be the four-manifold
$V_{p}=C_{0}^{p}\cup_{\phi_{0}\circ\sigma}Z$
where $C_{0}^{p}$ is the four-manifold obtained from $C_{0}$ by removing a neighbourhood of a simple closed curve and gluing back a copy of $S_{p}$. Moreover, we set
$\mathcal{E}^{\prime}_{p}=V^{\prime}\cup_{\zeta}V^{\prime}_{p}$
where $V^{\prime}=C_{0}\cup_{\phi_{0}}Z$ and $V^{\prime}_{p}=C_{0}^{p}\cup_{\phi_{0}}Z$ are made up by the same building blocks as $V$ and $V_{p}$ with different gluing maps. We refer the reader to [29, Construction 6.4] for further details.
The aim of this section is to provide a proof of the following proposition.
Proposition 14.
For every odd $p\in\mathbb{N}$, the following holds.
  1. (1)
    The four-manifold $\mathcal{E}^{\prime}_{p}$ is diffeomorphic to
    $X_{p}=(({\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})\setminus\nu(\gamma))\cup S_{p}=({\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})\mathbin{\#}\Sigma_{p},$
    where $\gamma\subset{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ is a simple loop.
  2. (2)
    The four-manifold $\mathcal{E}_{p}$ is homeomorphic but not diffeomorphic to $X_{p}$.
We proceed to provide some technical lemmas that are needed to distinguish the smooth structures of Theorem C. These results are generalizations of [29, Proposition 6.9] and [29, Lemma 6.3] respectively. In particular, we make use of the Ozsváth-Szabó four-manifold invariants defined in [31, Section 2.4].
In the following, we denote by $\Psi_{W,\mathfrak{t}}\in HF^{-}_{\text{red}}(\partial W,\mathfrak{t}|_{\partial W})$ the relative invariant of the $\operatorname{Spin}^{\mathbb{C}}$-structure $(W,\mathfrak{t})$ where $W$ is a compact four-manifold with non-empty boundary, and $\Phi_{X,\mathfrak{s},L}$ is the closed four-manifold invariant of the $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{s})$ with respect to a one-dimensional subspace $L\subset H_{2}(X;\mathbb{Q})$. We refer the reader to [29, Section 6.1] for the needed background on such invariants.
Lemma 15.
For any $p\geq 1$, there is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(V_{p},\mathfrak{t}_{p})$ such that $\Psi_{V_{p},\mathfrak{t}_{p}}\neq 0$. Moreover, all such ${\operatorname{Spin}}^{\mathbb{C}}$-structures can be chosen to coincide on $Z\subset V_{p}$. In particular, $V_{p}$ is homeomorphic but not diffeomorphic to $X_{0}(Q_{0})_{p}^{\prime}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ for
$X_{0}(Q_{0})_{p}^{\prime}=(X_{0}(Q_{0})\setminus\nu(\gamma))\cup S_{p},$
where $X_{0}(Q_{0})$ is the 00-trace of the Stevedore knot and $\gamma\subset X_{0}(Q_{0})$ is a null-homotopic simple loop.
[Proof]The case $p=1$ is [29, Proposition 6.9]. Since $C_{0}^{p}\setminus B^{4}$ is a rational homology cobordism between $S^{3}$ and $\partial C_{0}^{p}$, the same proof holds for $p>1$ by replacing $C_{0}$ with $C_{0}^{p}$ in the argument. The proof of the following result is a straightforward adaptation of the one of [29, Lemma 6.3].
Lemma 16.
Let $K$ be a fibered amphichiral knot of genus $g$. Let $W_{1}$ and $W_{2}$ be compact oriented 4-manifolds with positive second Betti number and boundary $S^{3}_{0}(K)$. Suppose there are ${\operatorname{Spin}}^{\mathbb{C}}$ structures $\{(W_{i},\mathfrak{u}_{i}):i=1,2\}$ such that $c_{1}(\mathfrak{u}_{i})$ evaluates to $22g-$ 2 on a generator of $H_{2}(S^{3}_{0}(K))$ and for which $\Psi_{W_{i},\mathfrak{u}_{i}}\neq 0$. Let $X$ be the result of gluing together $W_{1}$ and $W_{2}$ via any choice of orientation-reversing diffeomorphism of $S^{3}_{0}(K)$, and let $L\subset H_{2}(X;\mathbb{Q})$ be the line corresponding to the inclusion
$H_{2}(S^{3}_{0}(K))\rightarrow H_{2}(X).$
Then there is a ${\operatorname{Spin}}^{\mathbb{C}}$-structure $(X,\mathfrak{t})$ for which $\Phi_{X,\mathfrak{t},L}\neq 0$.
[Proof of Proposition 14] Item (1) follows from the construction and the fact that $\mathcal{E}^{\prime}$ is an exotic ${\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ by [29, Theorem 6.7]. By Lemma 7, there is a degree-two covering
$X_{p}\cong\mathcal{E}^{\prime}_{p}\xrightarrow{2:1}\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}.$
Let us now prove Item (2). In particular, $\mathcal{E}_{p}$ and $\mathcal{E}^{\prime}_{p}$ are homeomorphic by construction, since $\mathcal{E}$ and $\mathcal{E}^{\prime}$ are homeomorphic by [29, Theorem 6.7]. To conclude the proof, we need to show that the degree-two coverings $\mathcal{E}_{p}\rightarrow\mathcal{R}_{2p}$ and $\mathcal{E}^{\prime}_{p}\rightarrow\Sigma_{2}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ are not diffeomorphic. Consider the decomposition
(3.5) $\mathcal{E}_{p}=V_{a}\cup T_{Q_{0}}\cup V_{b}^{p}$
where $V_{a}$ and $V_{b}^{p}$ are copies of $V$ and $V_{p}$ respectively without the $T^{-}_{Q_{0}}$ component, see [29, Construction 6.4]. As for the case of $\mathcal{E}$, we see that $H_{2}(\mathcal{E}_{p})$ has a basis
$(E_{1},\dots,E_{4},E_{5}^{p},\dots,E_{8}^{p},\Xi,\Theta)$
where:
  • $(E_{1},\dots,E_{4})$ is a diagonal basis for $V_{a}$:
  • $(E_{5}^{p},\dots,E_{8}^{p}$ is a diagonal basis for $V_{b}^{p}$;
  • $(\Xi,\Theta)$ is the basis for $H_{2}(T_{Q_{0}})$ from [29, Remark 3.11]. In particular, $\Xi$ is represented by a capped-off Seifert surface for $Q_{0}$ in $S^{3}(Q_{0})$ and $\Theta$ is represented by a $2$-sphere of self-intersection $-2$ and $\Xi\cdot\Theta=1$.
As a consequence of Lemma 15 and 16, there is a ${\operatorname{Spin}}^{\mathbb{C}}$-structure $(\mathcal{E}_{p},\mathfrak{t})$ on such that $\Phi_{\mathcal{E}_{p},\mathfrak{t},L}\neq 0$, where $L=\text{Span}(\Xi)\subset H_{2}(\mathcal{E}_{p};\mathbb{Q})$. By [29, Lemma 6.2], we have that $\Xi$ cannot be represented by any surface of genus less than $2$. It is possible to show that if there was a diffeomorphism
$\mathcal{E}_{p}\rightarrow X_{p}$
then one could build a spherical representative for $\Xi$ as in the proof of [29, Theorem 6.7], which would lead to a contradiction.
Remark 17.
When $p$ is even, the four-manifold $D_{p}$ of Lemma 2 has two boundary components and the double covering $\mathcal{E}_{p}$ of $\mathcal{R}_{2p}$ is
$\mathcal{E}_{p}=(\mathcal{R}\setminus\nu(\alpha_{\mathcal{R}}))\cup D_{p}\cup(\mathcal{R}\setminus\nu(\alpha_{\mathcal{R}})).$
In particular, we have not been able to show that $\mathcal{R}_{2p}$ is exotic by reducing to the study of the Heegaard Floer invariants of $\mathcal{E}_{p}$ as in [29].
3.3. Proof of Theorem C
The four-manifold $\mathcal{R}_{2p}$ was constructed in (3.1) for any odd $p\in\mathbb{Z}_{>0}$ and its homeomorphism type was pinned down in Lemma 11; see Remark 12. The claim that there does not exist a diffeomorphism $\mathcal{R}_{2p}\rightarrow\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ follows from Proposition 14. Indeed, if any such diffeomorphism between $\mathcal{R}_{2p}$ and $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ were to exist, it would induce a diffeomorphism between the degree-two coverings
$\mathcal{E}_{p}\rightarrow(({\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})\setminus\nu(\gamma))\cup S_{p}=({\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})\mathbin{\#}\Sigma_{p}$
by Lemma 8 and Lemma 9. However, Proposition 14 indicates that these four-manifolds are not diffeomorphic. $\square$
4. Mechanism in Theorem D.
4.1. Seiberg-Witten invariants of smoothly reducible double coverings
The following proposition is a major trait in our proof of Theorem D. We direct the reader towards [34, Section 2.4], [8, Section 2.5] and the references there for the necessary background on Seiberg-Witten theory.
Proposition 18.
(Kotschick-Morgan-Taubes [27, Proposition 2], Baykur-Stipsicz-Szabó [8, Corollary 2.11]). Let $X_{1}$ and $X_{2}$ be closed oriented smooth four-manifolds and suppose that $b_{1}(X_{2})=0=b_{2}^{+}(X_{2})$. There is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{s}_{X})$ on the connected sum $X=X_{1}\mathbin{\#}X_{2}$ that extends a $\operatorname{Spin}^{\mathbb{C}}$-structure $(X_{1},\mathfrak{s}_{X_{1}})$ such that
$\operatorname{SW}_{X}(\mathfrak{s}_{X})=\pm\operatorname{SW}_{X_{1}}(\mathfrak{s}_{X_{1}}).$
In particular, if $X_{1}$ has nontrivial Seiberg-Witten invariant, then so does $X$.
Baykur-Stipsicz-Szabó observed in [8, Corollary 2.11] that Proposition 18 is a special case of the Seiberg-Witten blow-up formula [13]. Proposition 19, along with the blow up formula [13] (see [8, Proposition 2.8]). Proposition 18 will be used to argue the conclusion of the irreduciblity property in the statement of Theorem D, together with an argument due to Baykur-Stipsicz-Szabó [8].
Proposition 19.
(cf. [8, Proof of Proposition 6.2]). Let $\widetilde{M}$ be a closed smooth oriented simply connected four-manifold and assume it satisfies at least one of the following two properties.
$\bullet$ The four-manifold $\widetilde{M}$ has exactly two Seiberg-Witten basic classes ±$\pm K_{\widetilde{M}}\in\mathfrak{B}_{\widetilde{M}}$ or
$\bullet$ The four-manifold $\widetilde{M}$ is minimal and it has several Seiberg-Witten basic classes {±$\{\pm K^{1}_{\widetilde{M}},\ldots,\pm K^{i}_{\widetilde{M}}\}\in\mathfrak{B}_{\widetilde{M}}$ with $i\geq 2$.
Suppose there is an orientation-preserving fixed-point-free involution
$\tau:\widetilde{M}\rightarrow\widetilde{M}$
and denote by $M(2)=\widetilde{M}/\tau$ the quotient closed smooth oriented four-manifold. Let $\alpha_{M(2)}\subset M(2)$ a simple loop whose homotopy class generates the fundamental group $\pi_{1}(M(2))=\mathbb{Z}/2$.
The four-manifold
$M(2p)=(M(2)\setminus\nu(\alpha_{M(2)}))\cup S_{p}$
is irreducible, provided $p$ is an odd integer. Furthermore, there is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(\widehat{M}(p),\mathfrak{s}_{\widehat{M}(p)})$ on the degree-two covering
$\widehat{M}(p)\xrightarrow{2:1}M(2p)$
corresponding to the subgroup $\mathbb{Z}/p\subset\pi_{1}(M(2p))=\mathbb{Z}/2p$ for which the value of the Seiberg-Witten invariant satisfies
(4.1) $\operatorname{SW}_{\widehat{M}(p)}(\mathfrak{s}_{\widehat{M}(p)})=\operatorname{SW}_{\widetilde{M}\mathbin{\#}\Sigma_{p}}(\mathfrak{s_{\mathbin{\#}}})=\operatorname{SW}_{\widetilde{M}}(\mathfrak{s}_{\widetilde{M}})$
for $\operatorname{Spin}^{\mathbb{C}}$-structures $(\widetilde{M}\mathbin{\#}\Sigma_{p},\mathfrak{s}_{\mathbin{\#}})$ and $(\widetilde{M},\mathfrak{s}_{\widetilde{M}})$.
The reader might have already noticed that the properties that the simply connected four-manifold $\widetilde{M}$ of Proposition 19 is required to have imply that it is irreducible.
[Proof]Proceed by contradiction and suppose that there is a smooth connected sum decomposition
(4.2) $M(2p)=M_{1}\mathbin{\#}M_{2},$
where neither of the four-manifolds $M_{1}$ and $M_{2}$ is a homotopy four-sphere. Since the finite cyclic group $\pi_{1}(M(2p))=\mathbb{Z}/2p$ does not decompose as a free product of two non-trivial groups, without loss of generality we can assume that
$\pi_{1}(M_{1})=\mathbb{Z}/2p\quad\text{and}\quad\pi_{1}(M_{2})=\{1\}$
and $b_{2}(M_{2})\geq 1$. We hence obtain a decomposition of the degree-two covering
(4.3) $\widehat{M}(p)=\widehat{M}_{1}\mathbin{\#}2M_{2}$
from (4.2) where $\widehat{M}_{1}$ is the degree-two covering of $M_{1}$. Since by Lemma 8 there is an identification $\widehat{M}(p)=\widetilde{M}\mathbin{\#}\Sigma_{p}$, Proposition 18 implies that there is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(\widehat{M}(p),\mathfrak{s}_{\widehat{M}(p)})$ with $\operatorname{SW}_{\widehat{M}(p)}(\mathfrak{s}_{\widehat{M}(p)})\neq 0$ and it follows that $b_{2}^{+}(M_{2})=0$. In particular, the intersection form of $M_{2}$ is isomorphic to $b_{2}(M_{2})\langle-1\rangle.$
At this point, we employ an argument due to Baykur-Stipsicz-Szabo [8] to prove Proposition 19. Consider the map
(4.4) $c_{1}:\operatorname{Spin}^{\mathbb{C}}(\widehat{M}(p))\rightarrow H^{2}(\widehat{M}(p);\mathbb{Q})$
assigning to a given $\operatorname{Spin}^{\mathbb{C}}$-structure $(\widehat{M}(p),\mathfrak{s})$ its first Chern class with $\mathbb{Q}$-coefficients. Suppose first that the four-manifold $\widetilde{M}$ has exactly two Seiberg-Witten basic classes. As argued by Baykur-Stipsicz-Szabó [8, Proof of Proposition 6.2], the blow up formula [13] (see [8, Proposition 2.8]) and Proposition 18 imply that the image of the set
(4.5) $\{\mathfrak{s}_{\widehat{M}(p)}\in\operatorname{Spin}^{\mathbb{C}}(\widehat{M}(p)):\operatorname{SW}_{\widehat{M}(p)}(\mathfrak{s}_{\widehat{M}(p)})\neq 0\}$
under the map (4.4) has two elements, given that the four-manifold $\widetilde{M}$ is assumed to have exactly two Seiberg-Witten basic classes and the covering decomposes as $\widehat{M}(p)=\widetilde{M}\mathbin{\#}\Sigma_{p}$ by Lemma 8. On the other hand, the decomposition (4.3) implies that the image of the set (4.5) under (4.4) contains at least four elements. This is a contradiction and we conclude that the four-manifold $M(2p)$ is irreducible. A slight extension of Baykur-Stipsicz-Szabó’s argument allows us to reach a similar contradiction and conclude that $M(2p)$ is irreducible under the assumptions that the four-manifold $\widetilde{M}$ admits several Seiberg-Witten basic classes and that it is minimal. In this case, the image of the set (4.5) under the map (4.4) has $i$ elements, where $i>2$ is the number of Seiberg-Witten basic classes of $\widetilde{M}$. If there were indeed a decomposition as in (4.3), then the number of elements of the image of the set (4.5) under the map (4.4) would be strictly larger than $i$. We conclude that the four-manifold $M(2p)$ is irreducible. The claim on the value (4.1) of the Seiberg-Witten invariant is an immediate consequence of Proposition 18.
4.2. Proof of Theorem D
Denote by $X_{n}(2)\mathrel{\mathop{:}}=\widetilde{X}_{n}/\tau_{n}$ the closed smooth oriented four-manifold with fundamental group of order two that arises as the quotient of the closed oriented simply connected four-manifold $\widetilde{X}_{n}$ by its orientation-preserving fixed-point free involution $\tau_{n}:\widetilde{X}_{n}\rightarrow\widetilde{X}_{n}$. The classification result of Hambleton-Kreck [21, Theorem C] says that $X_{n}(2)$ is homeomorphic to $Y\mathbin{\#}\Sigma_{2,k}$ for a given closed simply connected topological four-manifold $Y$ with Kirby-Siebenmann invariant $\operatorname{KS}(Y)=k=\Sigma_{2,k}$. Fix an odd positive integer $p\in\mathbb{Z}_{>0}$ and assemble
(4.6) $X_{n}(2p)\mathrel{\mathop{:}}=(X_{n}(2)\setminus\nu(\alpha_{X_{n}(2)}))\cup S_{p},$
where $\alpha_{X_{n}(2)}\subset X_{n}(2)$ is a simple loop whose homotopy class is the generator
$\langle[\alpha_{(X_{n}(2)}]\rangle=\pi_{1}(X_{n}(2))=\mathbb{Z}/2$
and $S_{p}$ is the compact four-manifold that was manufactured in Section 2.1. Lemma 6 says that the fundamental group of the four-manifold (4.6) is the finite cyclic group $\pi_{1}(X_{n}(2p))=\mathbb{Z}/2p$. Moreover, there is a homeomorphism between $X_{n}(2p)$ and $Y\mathbin{\#}\Sigma_{2p,k}$ by Corollary 5 (and [21, Theorem C] as well).
Lemma 8 says that the degree-two covering $X_{n}(p)$ that corresponds to the subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p$ is diffeomorphic to $X_{n}(p)=\widetilde{X}_{n}\mathbin{\#}\Sigma_{p}$. Proposition 19 says that the four-manifold (4.6) is irreducible. At this point, we have proven all clauses of Theorem D except for the third one. To conclude on the existence of infinitely many pairwise non-diffeomorphic irreducible smooth structures on $Y\mathbin{\#}\Sigma_{2p,k}$ and conclude the proof of the theorem, we employ Lemma 9 and distinguish them in terms of their double coverings: the value of the Seiberg-Witten function is assumed to be $\operatorname{SW}_{\widetilde{X}_{n}}(\pm K_{n})=\pm n^{2}$ on its sole pair of basic classes and $\widetilde{X}_{n_{1}}$ is diffeomorphic to $\widetilde{X}_{n_{2}}$ if and only if $n_{1}=n_{2}$. Notice that this argument is due to Stipsicz-Szabó [34] and Baykur-Stipsicz-Szabó [8, Proposition 2.14]. $\square$
5. Proof of Theorem E: new irreducible examples with cyclic fundamental group.
Theorem D and the following result are the main ingredients in our proof of Theorem E.
Theorem 20.
(Arabadji-Morgan [4], Baykur-Stipsicz-Szabó [8], Stipsicz-Szabó [33, 34]). Let $(m,n)\in\mathbb{Z}_{>0}\times\mathbb{Z}_{>0}$ be positive integers that satisfy the identities
  • $4+5m\geq n$,
  • $4+5n\geq m$,
  • $m\neq 0\mod 2$ and
  • $m>7$ if $m=n$.
There is a closed smooth irreducible simply connected four-manifold $\widetilde{X}_{m,n}$ with
$b_{2}^{+}(\widetilde{X}_{m,n})=2m+1$ and $b_{2}^{-}(\widetilde{X}_{m,n})=2n+1$
that admits a fixed-point free orientation-preserving involution.
A substantial region of Theorem 20 is due to Arabadji-Morgan’s extensive study on the existence of fixed-point free orientation-preserving involutions on the closed irreducible symplectic simply connected four-manifolds constructed in [2, 19]. Theorem D follows from Theorem 20 and the last clause of Theorem D. $\square$
Remark 21.
As Porter Morgan kindly pointed out to us, several of the irreducible simply-connected four-manifolds with an involution constructed in [4] have more than two Seiberg-Witten basic classes. A stronger conclusion of infinitely many exotic smooth structures in Theorem E using our mechanism would depend on an improvement of Arabadji-Morgan’s result in this direction.
Remark 22.
It is interesting to understand what is obtained from applying our mechanism to other constructions of exotic four-manifolds like the recent ones due to Davis-Hayden-Huang-Ruberman-Sunukjian [11] and to Harris-Naylor-Park [23], for example.
6. Constructions of nonorientable four-manifolds and the examples of Theorem F.
6.1. Cut-and-paste procedure: nonorientable setting
There are two ways to lay down the procedure described in Section 2.1 in the nonorientable case. Notice that the framed simple loop $\alpha_{M}\subset M$ used in the construction (2.5) can be assumed to be orientation-preserving even if the ambient four-manifold $M$ is nonorientable and the construction described in that section results in a nonorientable four-manifold $M(p)$. We now describe the second version of this construction.
The nonorientable compact four-manifold $N_{p}$ of the second step of our procedure is constructed as follows; cf. Section 2.1. Denote the total space of the nonorientable three-sphere bundle over the circle by $S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ and let $\alpha\subset S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ be a simple loop whose homotopy class generates the infinite cyclic group $\pi_{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})=\mathbb{Z}$. Let $p$ be an odd integer and denote by $\alpha_{p}\subset S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ the orientation-reversing simple loop representing the class $[\alpha]^{p}\in H_{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1};\mathbb{Z})=\mathbb{Z}$. Its tubular neighborhood is diffeomorphic to the total space of the nonorientable three-disk bundle over the circle $\nu(\alpha_{p})=D^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$.
The compact nonorientable four-manifold
(6.1) $N_{p}=(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})\setminus\nu(\alpha_{p})$
with boundary $\partial N_{p}=S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ is the building block of the second step of our construction procedure when the goal is to produce nonorientable examples.
In the following, it will be useful to understand the universal cover of $N_{p}$, which we now describe via the following result.
Lemma 23.
For any odd $p\in\mathbb{Z}_{>0}$, the compact nonorientable four-manifold with boundary $N_{p}$ defined in (6.1) is double covered by the orientable four-manifold $S_{p}$ (2.1):
$S_{p}\xrightarrow{2:1}N_{p}.$
[Proof]The two four-manifolds $S_{p}$ and $N_{p}$ share the same universal cover. The deck transformations of $N_{p}$ are generated by the map
$\psi:\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})\rightarrow\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})$
given by
$(t,x)\mapsto(t+1,g_{p}(x)),$
where $g_{p}$ is the restriction of an orientation reversing automorphism of $S^{3}$ that cyclically permutes $B_{1},\dots,B_{p}$. The conclusion follows from the existence of a diffeomorphism
$S_{p}\rightarrow(\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i}))/\sim_{\psi^{2}}.$
Remark 24.
We can view $S^{3}$ by means of its genus one Heegaard splitting (2.3) and assume that the map $g_{p}$ in the proof of Lemma 23 is the composition of the map $f_{p}$ of Remark 3 and a reflection $\rho$ with respect to a two-sphere that intersects the first solid torus of (2.3) in $S^{1}\times D^{1}$ and the second one in the disjoint union of two $D^{2}$-fibers.
The following result records a key property of the four-manifold (6.1) that allows for the second and third step of our procedure to be successful in the nonorientable case. This property is an extension of a foundational result due to César de Sá [9] and Miller-Naylor [30], which states that any self-diffeomorphism of $S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ extends to a self-diffeomorphism of $D^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$.
Theorem 25.
(cf. Theorem A) For every odd integer $p\geq 1$ and for any self-diffeomorphism
$f:S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}\rightarrow S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1},$
there is a diffeomorphism
$F_{p}:N_{p}\rightarrow N_{p}$
such that $F_{p}|_{\partial}=f$.
Theorem A follows from Theorem 4 and Theorem 25.
[Proof]Kim-Raymond [24] (cf. [30]) showed that the diffeotopy group of $S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ is a quotient of the one of $S^{1}\times S^{2}$. This is isomorphic to $\mathbb{Z}/2\times\mathbb{Z}/2\times\mathbb{Z}/2$ and a set of generators is given by the Gluck twist (2.4) and the product maps $Id_{S^{1}}\times a$ and $r\times Id_{S^{2}}$, where $a$ is the antipodal involution of $S^{2}$ and $r$ is a orientation-reversing automorphism of $S^{1}$. We prove the extension result for the map $f:S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}\rightarrow S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ given by the Gluck twist and leave the cases of the other generators of the diffeotopy group of $S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ to the interested reader.
Without loss of generality, we can work in the context of Remark 24. Consider the map
$G:\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})\rightarrow S^{3}\setminus\sqcup_{i=1}^{p}B_{i}$
$(t,x)\mapsto r_{t}(x)$
where $r_{t}$ is a rotation of $S^{3}$ of angle $\pi t$, restricting to a rotation along the $D^{2}$-factor of $D^{2}\times S^{1}$ in (2.3). One can check that the smooth automorphism
$T:\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})\rightarrow\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})$
$(t,x)\mapsto(t,G(t,x))$
is $\psi$-equivariant (where $\psi$ is the map generating the deck transformations of $N_{p}$ in Lemma 23) and, therefore, it induces a smooth automorphism
$t_{p}:N_{p}\rightarrow N_{p}.$
In particular, the restriction of $t_{p}$ to the boundary is obtained by composing $p$ copies of the automorphism $t$. This yields a map isotopic to $t$, since $p$ was assumed to be odd.
Much like it was done in Section 2.1, remove the tubular neighborhood of an orientation-reversing simple loop $\alpha_{Y}\subset Y$ from a closed nonorientable four-manifold $Y$ to obtain a compact nonorientable four-manifold $Y\setminus\nu(\alpha_{Y})$ with boundary $\partial(Y\setminus\nu(\alpha_{Y}))=S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$. Cap it off with the compact four-manifold (6.1) to construct the closed nonorientable four-manifold
(6.2) $Y(p)\mathrel{\mathop{:}}=(Y\setminus\nu(\alpha_{Y}))\cup N_{p}.$
A useful consequence of Theorem 25 is recorded in the following corollary.
Corollary 26.
(cf. Theorem B). Let $Y_{1}$ and $Y_{2}$ be topological four-manifolds and let $\alpha_{Y_{i}}\subset Y_{i}$ be locally flat orientation-reversing simple loops, for $i=1,2$. Suppose there is a homeomorphism
$f:Y_{1}\setminus\nu(\alpha_{Y_{1}})\rightarrow Y_{2}\setminus\nu(\alpha_{Y_{2}}).$
The four-manifolds
$Y_{1}(p)\mathrel{\mathop{:}}=(Y_{1}\setminus\nu(\alpha_{Y_{1}}))\cup N_{p}\quad\text{and}\quad Y_{2}(p)\mathrel{\mathop{:}}=(Y_{2}\setminus\nu(\alpha_{Y_{2}}))\cup N_{p}$
are homeomorphic.
The simplicity of the construction makes it straight-forward to compute several basic invariants of the new four-manifold (6.2) as we now sample; see [22] for details on these invariants.
Lemma 27.
The closed nonorientable four-manifold $Y(p)$ defined in (6.2) has the following basic topological invariants for every $p\in\mathbb{Z}_{>0}$.
  1. (1)
    The Euler characteristic is $\chi(Y(p))=\chi(Y)$.
  2. (2)
    The second Stiefel-Whitney class is $w_{2}(Y(p))=w_{2}(Y)$ and the relation $w_{1}(Y(p))^{2}+w_{2}(Y(p))=w_{1}(Y)^{2}+w_{2}(Y)$ holds.
  3. (3)
    If there is a $\operatorname{Pin}^{\mathbb{C}}$-structure $(Y,\phi^{\mathbb{C}}_{Y})$, then there is a $\operatorname{Pin}^{\mathbb{C}}$-structure $(Y(p),\phi^{\mathbb{C}}_{Y(p)})$ and the Arf invariant is $\operatorname{arf}(Y,\phi^{\mathbb{C}}_{Y})=\operatorname{arf}(Y(p),\phi^{\mathbb{C}}_{Y(p)})$.
  4. (4)
    If the homotopy class of the simple loop $\alpha_{Y}\subset Y$ generates the cyclic fundamental group $\pi_{1}(Y)=\mathbb{Z}/q$, then the fundamental group of $Y(p)$ is $\pi_{1}(Y(p))=\langle t:t^{q\cdot p}=1\rangle$.
Much like the in the case of Lemma 6, a proof of Lemma 27 is obtained through straight-forward (co)homological computations, cobordism arguments and the Seifert-van Kampen theorem.
6.2. Nonorientable raw material and its properties
With the purpose of streamlining the exposition, we encapsule the background results we build on to prove Theorem F in the next triad of results. We first recall the building block that serves as raw material of the procedure we described in the introduction, along with several of its basic properties.
Theorem 28.
$\bullet$ Cappell-Shaneson [10]. There is a four-manifold $Q$ that is homeomorphic, but not diffeomorphic to ${\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}$. These two four-manifolds remain non-diffeomorphic under connect sums with arbitrarily many copies of $S^{2}\times S^{2}$.
$\bullet$ Gompf [19]. The universal covering $\widetilde{Q}$ is diffeomorphic to $S^{4}$.
$\bullet$ Akbulut [1, Theorem 2]. There is a diffeomorphism
$Q\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\rightarrow{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}.$
We now add a few words of the foundational results summarized in the previous theorem in order to put them into the context of our proof of Theorem F. Recall that the tubular neighborhood of
${\mathbb{R}\mkern-0.5mu\mathrm{P}}^{2}=\{[x:y:z:0:0]|x^{2}+y^{2}+z^{2}=1\}\subset S^{4}/{\sim}={\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}$
is the two-disk bundle over the real projective plane
$D^{2}\mathbin{\widetilde{\smash{\times}}}{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{2}=(D^{2}\times S^{2})/{\sim},$
where $(d,x)\sim(-d,\mathbb{A}(x))$ for every $d\in D^{2}$ and $x\in S^{2}$ and $\mathbb{A}:S^{2}\rightarrow S^{2}$ is the antipodal involution. The analysis done by Akbulut [1, Section 1] on the four-manifold $Q$ built by Cappell-Shaneson [10] reveals that the compact four-manifold $Z\mathrel{\mathop{:}}=Q\setminus\nu(\alpha_{Q})$ (where $\alpha_{Q}\subset Q$ is a simple loop that generates $\pi_{1}(Q)=\mathbb{Z}/2$) is homeomorphic, but not diffeomorphic to $D^{2}\mathbin{\widetilde{\smash{\times}}}{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{2}$. The second clause of Theorem 28 encapsules work of Gompf, where he showed that the orientation two-covering of $Z$ is diffeomorphic to $D^{2}\times S^{2}$. The third clause of Theorem 28 follows from the property $Z\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}=(D^{2}\mathbin{\widetilde{\smash{\times}}}{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{2})\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}$ [1, Section 3].
6.3. A spectral invariant as a diffeomorphism obstruction
The $\eta$-invariant will be used to distinguish the diffeomorphism type of our nonorientable four-manifolds. We refer the reader to [35] for further details on this spectral invariant and to [22, 26, 35] for background on $\operatorname{Pin}^{+}$-structures. We say that two four-manifolds $Y$ and $Y^{\prime}$ are stably diffeomorphic if there is a positive integer $k\in\mathbb{Z}_{>0}$ such that there is a diffeomorphism between the connected sums
$Y\mathbin{\#}(k-1)(S^{2}\times S^{2})$ and $Y^{\prime}\mathbin{\#}(k-1)(S^{2}\times S^{2})$,
where we set $0(S^{2}\times S^{2})=S^{4}$.
Theorem 29 (Stolz [35]).
$\bullet$ Two closed smooth nonorientable $\operatorname{Pin}^{+}$- four-manifolds $(Y,\phi_{Y})$ and $(Y^{\prime},\phi_{Y^{\prime}})$ with $H^{1}(Y;\mathbb{Z}/2)=\mathbb{Z}/2=H^{1}(Y^{\prime};\mathbb{Z}/2)$ are stably diffeomorphic if and only if their Euler characteristics agree and $\eta(Y,\phi_{Y})=\eta(Y^{\prime},\phi_{Y^{\prime}})\mod 2\mathbb{Z}$.
$\bullet$ There are $\operatorname{Pin}^{+}$-structures $({\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4},\pm\phi_{{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}})$ and $(Q,\pm\phi_{Q})$ such that
$\eta({\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4},\pm\phi_{{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}})=\pm\frac{1}{8}\mod 2\mathbb{Z}$
and
$\eta(Q,\pm\phi_{Q})=\pm\frac{7}{8}\mod 2\mathbb{Z}.$
In particular, $Q$ is not stably diffeomorphic to ${\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}$.
The last claim was proven by Cappell-Shaneson in [10].
6.4. $\operatorname{Pin}^{+}$-cobordism classes and a computational tool
We now elaborate on the role that the $\eta$-invariant plays in telling two given smooth structures apart. As we mentioned in the introduction, we distinguish the diffeomorphism classes by looking at the $\operatorname{Pin}^{+}$-structures on the tangent bundle of our four-manifolds through the $\eta$-invariant.
Theorem 30.
$\bullet$ Kirby-Taylor [26, Theorem 5.2]. There is a group isomorphism
$\Omega^{\operatorname{Pin}^{+}}_{4}\rightarrow\mathbb{Z}/16$
and the class of the real projective four-space
$[({\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4},\pm\phi_{{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}})]=\pm 1\in\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16.$
is a generator.
[…]
$\bullet$ Stolz [35].
$[(Q,\pm\phi_{Q})]=\pm 9\in\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16.$
We finish this section with the main computational tool to distinguish the four-manifolds of our procedure in the presence of a $\operatorname{Pin}^{+}$-structure.
Proposition 31.
Let $Y$ be a closed smooth four-manifold with first cohomology group $H^{1}(Y;\mathbb{Z}/2)=\mathbb{Z}/2$ and that admits a $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$. Let $\alpha_{Y}\subset Y$ be an orientation-reversing simple loop such that $[\alpha_{Y}]\neq 0\in H_{1}(Y;\mathbb{Z})$ and build
(6.3) $Y(p)\mathrel{\mathop{:}}=(Y\setminus\nu(\alpha_{Y}))\cup N_{p}.$
There are $\operatorname{Pin}^{+}$-structures
$(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1},\phi)$ and $(Y(p),\phi_{Y(p)})$
along with a $\operatorname{Pin}^{+}$-cobordism
(6.4) $((V,\phi_{V});(Y,\phi_{Y})\sqcup(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1},\phi),(Y(p),\phi_{Y(p)})).$
In particular, the identity
(6.5) $[(Y,\phi_{Y})]=[(Y(p),\phi_{Y(p)})]\in\Omega^{\operatorname{Pin}^{+}}_{4}$
holds.
A similar statement to Proposition 31 continues to hold if the simple loop $\alpha_{Y}\subset Y$ is assumed to be orientation-preserving.
[Proof]We first assemble the $\operatorname{Pin}^{+}$-structure $(Y(p),\phi_{Y(p)})$ by gluing together $\operatorname{Pin}^{+}$-structures $(Y\setminus\nu(\alpha_{Y}),\phi_{Y}^{0})$ and $(N_{p},\phi_{N_{p}})$ that induce the same $\operatorname{Pin}^{+}$-structure on their common boundary; see (6.3). The $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$ induces a $\operatorname{Pin}^{+}$-structure $(Y\setminus\nu(\alpha_{Y}),\phi_{Y}^{0})$ as a codimension zero submanifold [26, Construction 1.5], which restricts to a $\operatorname{Pin}^{+}$-structure
(6.6) $(\partial(Y\setminus\nu(\alpha_{Y})),\phi^{0}_{Y}|_{\partial})=(S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1},\phi_{\partial}).$
Since
$H^{1}(Y;\mathbb{Z}/2)=\mathbb{Z}/2=H^{1}(Y\setminus\nu(\alpha_{Y});\mathbb{Z}/2)=H^{1}(\partial(Y\setminus\nu(\alpha_{Y}));\mathbb{Z}/2),$
there are two choices of such structures [35, Corollary 6.4] and there is a one-to-one correspondence between the initial $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$ and the $\operatorname{Pin}^{+}$-structure (6.6) (see [25, Proposition 1 Chapter IV]). We now argue that there is a $\operatorname{Pin}^{+}$-structure $(N_{p},\phi_{N_{p}})$ that induces the $\operatorname{Pin}^{+}$-structure (6.6) on $\partial N_{p}=S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$. Since $w_{2}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})=0$ and $H^{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1};\mathbb{Z}/2)=\mathbb{Z}/2$, there are two $\operatorname{Pin}^{+}$-structures on $S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ [35, Corollary 6.4]. The hypothesis $p\neq 0\mod 2$ allows us to pick the $\operatorname{Pin}^{+}$-structure $(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1},\phi)$ that induces $(N_{p},\phi_{N_{p}})$ such that the induced $\operatorname{Pin}^{+}$-structure $(\partial N_{p},\phi_{N_{p}}|_{\partial})$ matches (6.6). We conclude that there is a $\operatorname{Pin}^{+}$-structure $(Y(p),\phi_{Y(p)})$. Next, we put together the compact five-manifold $V$ by gluing the product cobordisms
$Y\times[0,1]$ and $(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})\times[0,1]$
along the upper lids $Y\times\{1\}$ and $(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})\times\{1\}$ according to the construction of $Y(p)$ in (6.3). The $\operatorname{Pin}^{+}$-structure $(V,\phi_{V})$ is assembled analogously to the assembly of $(Y,\phi_{Y})$ that was described before. Given that
$[(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1},\pm\phi)]=0\in\Omega^{\operatorname{Pin}^{+}}_{4},$
the identity (6.5) follows from the existence of the cobordism (6.4).
6.5. Proof of Theorem F
Fix an odd $p\in\mathbb{Z}_{\geq 0}$ and consider the compact four-manifold $N_{p}$ that was defined in (6.1). Assemble the nonorientable four-manifolds
(6.7) $A(2p)\mathrel{\mathop{:}}=({\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}\setminus\nu(\alpha_{{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}}))\cup N_{p}=(D^{2}\mathbin{\widetilde{\smash{\times}}}{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{2})\cup N_{p}$
and
(6.8) $B(2p)\mathrel{\mathop{:}}=(Q\setminus\nu(\alpha_{Q}))\cup N_{p}=Z\cup N_{p},$
where $Q$ is Cappell-Shaneson’s exotic copy of the real projective four-space [10] discussed in Theorem 28. Lemma 27 implies that the four-manifolds (6.7) and (6.8) are $\mathbb{Q}$-homology real projective four-spaces whose fundamental group is a finite cyclic group of order $2p$. This establishes the claim of Item (1) of Theorem F. Moreover, the four-manifolds $A(2p)$ and $B(2p)$ are homeomorphic by Corollary 26 since $Q$ is homeomorphic to ${\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}$. This settles the claim of Item (2). Since there are $\operatorname{Pin}^{+}$-structures $({\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4},\pm\phi_{{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}})$ and $(Q,\pm\phi_{Q})$, Proposition 31 implies that the four-manifolds (6.7) and (6.8) admit a $\operatorname{Pin}^{+}$-structure. There are actually two such structures. Indeed, the first cohomology groups with $\mathbb{Z}/2$-coefficients of the four-manifolds (6.7) and (6.8) are
$H^{1}(A(2p);\mathbb{Z}/2)=\mathbb{Z}/2=H^{1}(B(2p);\mathbb{Z}/2)$
and it follows that there are two $\operatorname{Pin}^{+}$-structures $(A(2p),\pm\phi_{A(2p)})$ and two $\operatorname{Pin}^{+}$-structures $(B(2p),\pm\phi_{B(2p)})$ [35, Corollary 6.4], . Moreover, the $\operatorname{Pin}^{+}$-structures on the same underlying smooth four-manifold are mutual inverses in the fourth $\operatorname{Pin}^{+}$-cobordism group $\Omega^{\operatorname{Pin}^{+}}_{4}$ and the values
(6.9) $\eta(A(2p),\pm\phi_{A(2p)})=\pm\frac{1}{8}\mod 2\mathbb{Z}$
and
(6.10) $\eta(B(2p),\pm\phi_{B(2p)})=\pm\frac{7}{8}\mod 2\mathbb{Z}$
hold by Proposition 31, Theorem 29 and Theorem 30. Using Theorem 29, we conclude that the connected sums
$A(2p)\mathbin{\#}(k-1)(S^{2}\times S^{2})$ and $B(2p)\mathbin{\#}(k-1)(S^{2}\times S^{2})$
belong to different classes in $\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16$ for any $k\in\mathbb{Z}_{>0}$ and that the four-manifolds $A(2p)$ and $B(2p)$ are not stably-diffeomorphic. This settles the claim of Item (5). Item (6) follows from the third clause of Theorem 28 and the corresponding coverings of the decompositions (6.7) and (6.8).
Finally, to construct the closed non-smoothable topological four-manifold $\ast A(2p)$, we substitute $\mathbb{R}P^{4}$ with the closed non-smootable four-manifold $\ast\mathbb{R}P^{4}$ that is homotopy equivalent to the real projective four-space in (6.7). The latter four-manifold exists by a computation using the surgery exact sequence [39, Chapter 14] and an explicit construction of it was given by Ruberman in [32]. The existence and uniqueness results of a topological tubular neighborhood $\nu(\alpha_{\ast\mathbb{R}P^{4}})$ for a locally flat simple loop $\alpha_{\ast\mathbb{R}P^{4}}\subset\ast\mathbb{R}P^{4}$ are found in [16, Section 5.1]. Choose a loop that generates the fundamental group and assemble
(6.11) $\ast A(2p)\mathrel{\mathop{:}}=(\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}\setminus\nu(\alpha_{\ast{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}}))\cup N_{p}.$
The existence of a homotopy equivalence $\ast\mathbb{R}P^{4}\rightarrow\mathbb{R}P^{4}$ implies that there is a homotopy equivalence $\ast A(2p)\rightarrow A(2p)$. From (6.11) and the related cobordism, we have that the Kirby-Siebenmann invariant is $\operatorname{KS}(\ast A(2p))=\operatorname{KS}(\ast\mathbb{R}P^{4})=1\in\mathbb{Z}/2$ [16, Theorem 9.2] (see [16, Section 9.2] for a description of the construction of this invariant). Thus, the four-manifold $\ast A(2p)$ is non-smoothable for any odd $p\in\mathbb{Z}_{>0}$. $\square$